Solve by substitution
9x+4y/3=5
4x-y/3=5

Answer by Maths68(1474)

9x+4y/3=5 ............(1)
4x-y/3=5 .............(2)
Multiply (1) by 3 both sides of the equation
3*(9x+4y/3)=3*5
27x+4y=15.............(3)
Multiply (2) by 3 both sides of the equation
3*(4x-y/3)=3*5
12x-y=15...............(4)
Multiply (4) by 4 both sides and add to (3)
4*(12x-y)=4*15...............(4)
48x-4y=60
27x+4y=15
-------------
75x=75
Divide above equation by 75 both sides
75x/75=75/75
x=1
Put the value of x in (3)
27x+4y=15.............(3)
27(1)+4y=15
27+4y=15
4y=15-27
4y=-12
Divide by 4 both side of above equation
4y/4=-12/4
y=-3
(x,y)=(1, -3)
Check
Put the value of x and y in (1)
9x+4y/3=5 ............(1)
9(1)+4(-3)/3=5
9-4=5
5=5
Put the value of x and y in (2)
4x-y/3=5 .............(2)
4(1)-(-3)/3=5
4+1=5
5=5